# Learning with Statistical Equality Constraints

### v1

@@ 3 Generalization error
Notice that the Lagrangian for (Pϕ) is also defined as in (1). Similarly, its dual problem is given by
@@
$R(\nu)$ measures the maximum distance between $\Theta_{\nu}$ and $\text{Feas}(\text{P})$. Hence, it is a decreasing function of $\nu$ and vanishes when $\nu=0$. However, if relaxing the constraints causes the feasibility set to change radically, $R(\nu)$ could be unbounded for $\nu>0$. This can be avoided by ensuring that $\text{Feas}(\text{P}_{\theta\nu})$ is bounded, by e.g., regularizing $\Theta$ by enforcing $\left\|\theta\right\|_{2}\leq W$ for some $W>0$.
@@
for $i=1,\ldots,I$ and $j=1,\ldots,J$ (see [7, Equation 5.58]). Though differentiability and strong duality of (P-pert) does not generally hold in our setting, approximate relations can be obtained in terms of the bound on the duality gap $P^{\star}-D^{\star}$ (see Remark B.6). Thus, $\left\|\gamma^{\star}\right\|_{1}=\left\|\lambda^{\star}\right\|_{1}+\left\|\mu^{\star}\right\|_{1}$, that appears in (8), measures how much the objective changes when the constraints are perturbed. Constraint sensitivity also appears in terms of $R(\nu)$, which marks a clear difference with the generalization bounds for inequality constraints ($J=0$) that depend solely on $\left\|\lambda^{\star}\right\|_{1}$ (see Remark 3.1). The bound in (8) also depends on the sensitivity of the loss functions and the model (with respect to its parameters) through their Lipschitz constants $L$ and $L_{\theta}$ respectively.
@@
$\mathcal{M}_{\phi}=\left\{(f,\mathbf{u},\mathbf{v})\in\mathbb{R}^{I}\times\mathbb{R}^{J}\;\middle|\;\begin{aligned} \exists\phi\in\Phi\quad\text{ s.t.}\quad&\mathbb{E}_{\mathbb{D}}\left[\ell(\phi(\mathbf{x}),y)\right]=f,\\ &\mathbb{E}_{\mathbb{P}_{i}}\left[g_{i}(\phi(\mathbf{x}),y)\right]\leq u_{i},\quad\text{for }i=1,\ldots,I\text{,}\\ \text{ and }\quad&\mathbb{E}_{\mathbb{Q}_{j}}\left[h_{j}(\phi(\mathbf{x}),y)\right]=v_{j}\quad\text{for }j=1,\ldots,J\text{,}\end{aligned}\right\}$
@@ 4 Algorithm
Suppose Assumptions 1 and 5 hold and let the loss functions $\ell,g_{i},h_{j}$ be $B$-bounded. Let $U_{0}=\underset{\gamma^{\star}\in\text{Opt}(\hat{\text{D}})}{\text{inf}}\;\left\|\gamma^{(0)}-\gamma^{\star}\right\|$. Then, for any $T\in\mathbb{N}$, it holds that, $\left|\hat{D}^{\star}-\frac{1}{T}\sum_{t=0}^{T-1}\left(\hat{L}(\theta^{(t)},\gamma^{(t)})\right)\right|\leq\rho+\frac{U_{0}}{2\eta T}+\frac{1}{2}(I+J)\eta B^{2}.$ (10) If $\eta\leq\frac{\rho}{(I+J)B^{2}}$ and $T\geq\frac{U_{0}}{\eta\rho}$, then the bound is equal to $2\rho$.
@@ B.1.1 Bochner spaces
If both limits on the RHS exist, then the limit on the LHS exists as well ([51] Theorem 3.4). This proves that $f\in L^{1}(\mathbb{B})\cap L^{1}(\mathbb{P})\Rightarrow f\in L^{1}(\mathbb{B}+\mathbb{P})$.
@@ B.1.3 Constrained optimization problems
Let $\text{Feas}(\text{P}_{0})$ refer to the subset of $\mathcal{X}$ that satisfies all the constraints, i.e. the set of feasible solutions. If ($\text{P}_{0}$) is infeasible, we set $P_{0}^{\star}$ to $\infty$. If $P_{0}^{\star}=$-$\infty$ then we say that ($\text{P}_{0}$) is unbounded. If $\ell$ is bounded, then ($\text{P}_{0}$) is also bounded (from below). Let $\text{Opt}(\text{P}_{0})$ refer to the subset of $\text{Feas}(\text{P}_{0})$ that achieves $P_{0}^{\star}$, i.e. the set of optimal solutions. Closely related to ($\text{P}_{0}$) is its Lagrangian function,
@@
and its dual function is $q_{0}(\lambda,\mu)=\underset{x\in\mathcal{X}}{\text{inf}}\;L_{0}(x,\lambda,\mu)$. The dual function defines the dual problem,
@@
We call $\lambda\in\mathbb{R}_{\geq 0}^{I},\mu\in\mathbb{R}^{J}$ the dual variables and denote optimal dual variables, when they exist, with a $\star$, such as $(\lambda^{\star},\mu^{\star})$. We denote the set of all optimizers of ($\text{D}_{0}$) as $\text{Opt}(\text{D}_{0})$, which are also often called Lagrange multipliers. Let $F_{0}:\mathcal{X}\to\mathbb{R}^{1+I+J}$ be the vector function obtained by stacking the objective and the constraint functions, i.e.,
@@
$\mathcal{M}_{0}=F_{0}(\mathcal{X})+\mathbb{R}_{\geq 0}^{1+I}\times\left\{0\right\}^{J}.$
@@
$\mathcal{C}_{0}$ $=\left\{(g,h)\in\mathbb{R}^{I+J}\;|\;(\ell,g,h)\in\mathcal{M}_{0}\right\}$
$=C_{0}(\mathcal{X})+\mathbb{R}_{\geq 0}^{I}\times\left\{0\right\}^{J}.$ (24)
where $C_{0}(x)=\left[g_{1}(x),\ldots g_{I}(x),h_{1}(x),\ldots h_{J}(x)\right]^{\top}$ is the vector function formed by stacking only the $I+J$ constraint functions. We call $C_{0}$ the constraint vector and $\mathcal{C}_{0}$ the constraint epigraph. The relative interior of $\mathcal{C}_{0}$ is defined as,
@@ B.1.4 Geometric conditions for strong duality
$P_{0}^{\star}=\underset{(\ell,0^{I},0^{J})\in\mathcal{M}_{0}}{\text{inf}}\;\ell.$ (25)
@@
$D_{0}^{\star}=\underset{(\ell,0^{I},0^{J})\in\overline{\text{conv}}\left(\mathcal{M}_{0}\right)}{\text{inf}}\;\ell$. (26)
Equations (25) and (26) suggest that the convexity and closure of $\mathcal{M}_{0}$ is closely related to strong duality. In fact, convexity and closure are together sufficient for strong duality [5, Proposition 4.3.2]. However, the requirement for closure can also be dropped when the target constraint values ($0^{I+J}$) are in the relative interior of the constraint set $\mathcal{C}_{0}$, as seen in Theorem B.1. Theorem B.1 also provides sufficient conditions (i.e. the non-extremality of the constraint levels) for the existence of Lagrange multipliers. Note that this is non-trivial, since strong duality is not always sufficient for the existence of Lagrange multipliers, for example, see [5, Example 5.3.3] where $\mathcal{M}_{0}$ is both convex and closed, but $\text{Opt}(\text{D}_{0})=\emptyset$.
@@ B.1.5 Constraint perturbations
Upper bound . Let $\lambda\in\mathbb{R}_{\geq 0}^{I}$ and $\mu\in\mathbb{R}^{J}$ be arbitrarily chosen. Let $\mu_{+}$ and $\mu_{-}$ be defined pointwise as $\mu_{+,j}=\max\left\{0,\mu_{j}\right\}$ and $\mu_{-,j}=\max\left\{0,-\mu_{j}\right\}$. Clearly $\mu=\mu_{+}-\mu_{-}$ and $\left\|\mu_{+}\right\|_{1}+\left\|\mu_{-}\right\|_{1}=\left\|\mu\right\|_{1}$. Applying these facts, we can verify that $L_{0}$ can be rewritten as follows,
@@
Lower bound . Now let us prove the lower bound. Let $\lambda\in\mathbb{R}_{\geq 0}^{I}$ and $\mu_{+},\mu_{-}\in\mathbb{R}_{\geq 0}^{J}$ be arbitrarily chosen. Then we can rewrite $L_{\epsilon}$ as follows,
@@ B.2.1 Convexity of $\mathcal{M}_{\phi}$
Recall that $\mathcal{M}_{\phi}=F_{\phi}(\Phi)+\mathbb{R}_{\geq 0}^{1+I}\times\mathbf{0}^{J}$. Clearly $\mathbb{R}_{\geq 0}^{1+I}\times\mathbf{0}^{J}$ is a convex set, and since the Minkowski sum of two convex sets is convex ([50] Theorem 3.1), therefore $\mathcal{M}_{\phi}$ is also convex, which concludes the proof. ∎
@@ B.2.2 Strong duality of (Pϕ)
Assumption 2 The premise of the proposition, along with atomlessness from Assumption 2 allows us to invoke Lemma B.5 to conclude that $\mathcal{M}_{\phi}$ is a convex set.
@@ B.3 Duality gap
Note that the set $\Theta_{\nu}$ (defined in Section 3) is nothing but the feasibility set of ($\text{P}_{\theta\nu}$), i.e. $\text{Feas}(\text{P}_{\theta\nu})$. Let its dual value be $D^{\star}_{\theta\nu}$, defined similarly to $D^{\star}$ (Section 3). We define its functional version,
@@
Let its dual value be $D^{\star}_{\phi\nu}$. We will bound the duality gap based on (5) which uses the functional strong duality relation $P^{\star}_{\phi}=D^{\star}_{\phi}$, and which we further decompose as,
@@ B.3.4 Duality gap of ($\text{P}_{\theta\nu}$)
Let $\gamma\in\mathbb{R}_{\geq 0}^{I+2J}$ denote the dual variable for ($\text{P}_{\theta\nu}$) or ($\text{P}_{\phi\nu}$). Clearly $L_{\theta\nu}(\theta,\gamma)=L_{\phi\nu}(f_{\theta},\gamma)$. Therefore,
$q_{\theta\nu}(\gamma)=\underset{\theta\in\Theta}{\text{inf}}\;L_{\theta\nu}(\theta,\gamma)$ $=\underset{f_{\theta}\in\mathcal{H}}{\text{inf}}\;L_{\phi\nu}(f_{\theta},\gamma)\geq\underset{\phi\in\Phi}{\text{inf}}\;L_{\phi\nu}(\phi,\gamma)=q_{\phi\nu}(\gamma).$
@@
$D^{\star}_{\theta\nu}=\sup_{\gamma\in\mathbb{R}_{\geq 0}^{I+2J}}q_{\theta\nu}(\gamma)\geq\sup_{\gamma\in\mathbb{R}_{\geq 0}^{I+2J}}q_{\phi\nu}(\gamma)=D^{\star}_{\phi\nu}$
@@ B.4 Dual estimation error
$D^{\star}-\hat{D}^{\star}$ $=q(\gamma^{\star})-\hat{q}(\hat{\gamma}^{\star})$
$\leq q(\gamma^{\star})-\hat{q}(\gamma^{\star}),$ (49)
where the inequality follows from the suboptimality of $\gamma^{\star}$ with respect to ($\hat{\text{D}}$), i.e. because $\hat{q}(\gamma^{\star})\leq\hat{q}(\hat{\gamma}^{\star})=\hat{D}^{\star}$. Let $\epsilon>0$ be arbitrary and let $\theta$ be $\epsilon$-optimal with respect to $\hat{q}(\gamma^{\star})$, i.e,
$\hat{q}(\gamma^{\star})\leq\hat{L}(f_{\theta},\gamma^{\star})\leq\hat{q}(\gamma^{\star})+\epsilon.$ (50)
Clearly, $q(\gamma^{\star})\leq L(f_{\theta},\gamma^{\star})$ since $q(\gamma^{\star})=\underset{\theta}{\text{inf}}\;L(f_{\theta},\gamma^{\star})$. This fact and (50), applied to (49), implies that,
$D^{\star}-\hat{D}^{\star}\leq q(\gamma^{\star})-\hat{q}(\gamma^{\star})\leq L(f_{\theta},\gamma^{\star})-\hat{L}(f_{\theta},\gamma^{\star})+\epsilon.$ (51)
Now let us expand the difference of Lagrangians in the RHS of (51), also splitting $\gamma^{\star}$ into its components $(\lambda^{\star},\mu^{\star})$,
@@
This time we pick some $\theta$ that is $\epsilon$-optimal with respect to $q(\hat{\gamma}^{\star})$ (instead of $\hat{q}(\gamma^{\star})$). Since (53) holds for any $\theta$ (with probability greater than $1-(1+I+J)\delta$) therefore Equations (55) and (56) hold simultaneously with probability greater than the same bound, $1-(1+I+J)\delta$. Therefore, they together imply that $|D^{\star}-\hat{D}^{\star}|\leq(1+\max{}\left\{\left\|\gamma^{\star}\right\|_{1},\left\|\hat{\gamma}^{\star}\right\|_{1}\right\})\zeta^{\text{UC}}(N_{\text{min}},\delta)$, which completes the proof. ∎
@@
$\mathcal{M}=\left\{(f,\mathbf{u},\mathbf{v})\in\mathbb{R}^{I}\times\mathbb{R}^{J}\;\middle|\;\begin{aligned} \exists\theta\in\Theta\quad\text{ s.t.}\quad&\mathbb{E}_{\mathbb{D}}\left[\ell(f_{\theta}(\mathbf{x}),y)\right]=f,\\ &\mathbb{E}_{\mathbb{P}_{i}}\left[g_{i}(f_{\theta}(\mathbf{x}),y)\right]\leq u_{i},\quad\text{for }i=1,\ldots,I\text{,}\\ \text{ and }\quad&\mathbb{E}_{\mathbb{Q}_{j}}\left[h_{j}(f_{\theta}(\mathbf{x}),y)\right]=v_{j}\quad\text{for }j=1,\ldots,J\end{aligned}\right\}.$
@@
In this section we will provide guarantees for Algorithm 1, which approximately solves ($\hat{\text{D}}$). We start by defining the empirical dual function in terms of the empirical Lagrangian (2),
$\hat{q}(\lambda,\mu)=\underset{\theta\in\Theta}{\text{inf}}\;\hat{L}(f_{\theta},\lambda,\mu).$
The empirical dual value can be written as $\hat{D}^{\star}=\sup_{\lambda\in\mathbb{R}^{I}_{+},\,\mu\in\mathbb{R}^{J}}\hat{q}(\lambda,\mu)$. Let $\gamma$ denote the tuple $(\lambda,\mu)$ in the sequel, and similarly for $\gamma^{\star},\hat{\gamma}^{\star},\gamma^{(t-1)}$ etc. We can prove that the set of supergradients for $\hat{q}(\gamma)$ at $\gamma$ is equal to the convex hull $\text{conv}\left(\left\{\hat{C}(\theta)\;|\;\theta\in\Theta,L(f_{\theta},\gamma)=\hat{q}(\gamma)\right\}\right)$ (e.g. with [53, Theorem 2.87]). We will now show that approximate minimizers produce approximate supergradients. Recall that $\hat{F}$ is the cost-constraint vector of the empirical problem.
@@
Let $\theta^{(t)}$ satisfy $\hat{q}(\gamma^{(t)})\leq\hat{L}(f_{\theta},\gamma^{(t)})\leq\hat{q}(\gamma^{(t)})+\rho$. Then, it follows that, for all $\gamma^{\prime}\in\mathbb{R}_{\geq 0}^{I}\times\mathbb{R}^{J}$
$\hat{q}(\gamma^{(t)})\geq\hat{q}(\gamma^{\prime})+\sum_{k=1}^{K}(\gamma^{(t)}_{k}-\gamma^{\prime}_{k})\hat{F}_{k}(\theta^{(t)})-\rho.$
@@
By the definition of the dual function, we have for any $\gamma^{\prime}\in\mathbb{R}_{\geq 0}^{I}\times\mathbb{R}^{J}$,

### v2

@@ 3 Generalization error
Notice that the Lagrangian for (Pϕ) is the same as in (1). Similarly, its dual problem is given by
@@
$R(\nu)$ measures the maximum distance between $\Theta_{\nu}$ and $\text{Feas}(\text{P})$. Hence, it is a decreasing function of $\nu$ and vanishes when $\nu=0$. However, if relaxing the constraints causes the feasibility set to change radically, $R(\nu)$ could be unbounded for $\nu>0$. This can be avoided by ensuring that $\Theta_{\nu}$ is bounded, by e.g., regularizing $\Theta$ by enforcing $\left\|\theta\right\|_{2}\leq W$ for some $W>0$.
@@
for $i=1,\ldots,I$ and $j=1,\ldots,J$ (see [7, Equation 5.58]). Though differentiability and strong duality of (P-pert) does not generally hold in our setting, approximate relations can be obtained in terms of the bound on the duality gap $P^{\star}-D^{\star}$ (see Remark B.6). Thus, $\left\|\gamma^{\star}\right\|_{1}=\left\|\lambda^{\star}\right\|_{1}+\left\|\mu^{\star}\right\|_{1}$, that appears in (8), measures how much the objective changes when the constraints are perturbed. Constraint sensitivity also appears in terms of $R(\nu)$, which marks a clear difference from the generalization bounds for inequality constraints ($J=0$) that depend solely on $\left\|\lambda^{\star}\right\|_{1}$ (see Remark 3.1). The bound in (8) also depends on the sensitivity of the loss functions and the model (with respect to its parameters) through their Lipschitz constants $L$ and $L_{\theta}$ respectively.
@@
$\mathcal{M}_{\phi}=\left\{(f,\mathbf{u},\mathbf{v})\in\mathbb{R}^{1+I+J}\;\middle|\;\begin{aligned} \exists\phi\in\Phi\quad\text{ s.t.}\quad&\mathbb{E}_{\mathbb{D}}\left[\ell(\phi(\mathbf{x}),y)\right]=f,\\ &\mathbb{E}_{\mathbb{P}_{i}}\left[g_{i}(\phi(\mathbf{x}),y)\right]\leq u_{i},\quad\text{for }i=1,\ldots,I\text{,}\\ \text{ and }\quad&\mathbb{E}_{\mathbb{Q}_{j}}\left[h_{j}(\phi(\mathbf{x}),y)\right]=v_{j}\quad\text{for }j=1,\ldots,J\text{,}\end{aligned}\right\}$
@@ 4 Algorithm
Suppose Assumptions 1 and 5 hold and let the loss functions $\ell,g_{i},h_{j}$ be $B$-bounded. Let $U_{0}=\underset{\gamma^{\star}\in\text{Opt}(\hat{\text{D}})}{\text{inf}}\;\left\|\gamma^{(0)}-\gamma^{\star}\right\|$. Then, for any $T\in\mathbb{N}$, it holds that, $\left|\hat{D}^{\star}-\frac{1}{T}\sum_{t=0}^{T-1}\left(\hat{L}(f_{\theta^{(t)}},\gamma^{(t)})\right)\right|\leq\rho+\frac{U_{0}}{2\eta T}+\frac{1}{2}(I+J)\eta B^{2}.$ (10) If $\eta\leq\frac{\rho}{(I+J)B^{2}}$ and $T\geq\frac{U_{0}}{\eta\rho}$, then the bound is equal to $2\rho$.
@@ B.1.1 Bochner spaces
If the limits (with respect to $n$) of both terms on the RHS exist, then the limit on the LHS exists as well ([51] Theorem 3.4). This proves that $f\in L^{1}(\mathbb{B})\cap L^{1}(\mathbb{P})\Rightarrow f\in L^{1}(\mathbb{B}+\mathbb{P})$.
@@ B.1.3 Constrained optimization problems
Let $\text{Feas}(\text{P}_{0})$ refer to the subset of $\mathcal{X}$ that satisfies all the constraints, i.e. the set of feasible solutions. If ($\text{P}_{0}$) is infeasible, we set $P_{0}^{\star}$ to $\infty$. If $P_{0}^{\star}=$-$\infty$ then we say that ($\text{P}_{0}$) is unbounded. If $\ell$ is bounded, then ($\text{P}_{0}$) is also bounded (from below). Let $\text{Opt}(\text{P}_{0})$ refer to the subset of $\text{Feas}(\text{P}_{0})$ that achieves $P_{0}^{\star}$, i.e. the set of optimal solutions. Closely related to ($\text{P}_{0}$) is its Lagrangian,
@@
where $\lambda\in\mathbb{R}_{+}^{I}$ (non-negative) and $\mu\in\mathbb{R}^{J}$ are the dual variables, collecting $\lambda_{i}$ and $\mu_{j}$ respectively. The dual function of ($\text{P}_{0}$) is $q_{0}(\lambda,\mu)=\underset{x\in\mathcal{X}}{\text{inf}}\;L_{0}(x,\lambda,\mu)$. The dual function defines the dual problem,
@@
The optimal dual variables, when they exist, are denoted with a $\star$, such as $(\lambda^{\star},\mu^{\star})$. We denote the set of all optimizers of ($\text{D}_{0}$) as $\text{Opt}(\text{D}_{0})$, which are also often called Lagrange multipliers. Let $F_{0}:\mathcal{X}\to\mathbb{R}^{1+I+J}$ be the vector function obtained by stacking the objective and the constraint functions, i.e.,
@@
$\mathcal{M}_{0}=F_{0}(\mathcal{X})+\mathbb{R}_{+}^{1+I}\times\left\{0\right\}^{J}.$
@@
$\mathcal{C}_{0}$ $=\left\{(u,v)\in\mathbb{R}^{I+J}\;|\;(f,u,v)\in\mathcal{M}_{0}\right\}$
$=C_{0}(\mathcal{X})+\mathbb{R}_{+}^{I}\times\left\{0\right\}^{J}.$ (24)
where $C_{0}(x)=\left[g_{1}(x),\ldots,g_{I}(x),h_{1}(x),\ldots,h_{J}(x)\right]^{\top}$ is the vector function formed by stacking only the $I+J$ constraint functions. We call $C_{0}$ the constraint vector and $\mathcal{C}_{0}$ the constraint value epigraph. The relative interior of $\mathcal{C}_{0}$ is defined as,
@@ B.1.4 Geometric conditions for strong duality
$P_{0}^{\star}=\underset{(f,0^{I},0^{J})\in\mathcal{M}_{0}}{\text{inf}}\;f.$ (25)
@@
$D_{0}^{\star}=\underset{(f,0^{I},0^{J})\in\overline{\text{conv}}\left(\mathcal{M}_{0}\right)}{\text{inf}}\;f.$ (26)
Equations (25) and (26) suggest that the convexity and closure of $\mathcal{M}_{0}$ is closely related to strong duality. In fact, convexity and closure are together sufficient for strong duality [5, Proposition 4.3.2]. However, the requirement for closure can also be dropped when the target value of the constraints ($0^{I+J}$) is in the relative interior of the constraint set $\mathcal{C}_{0}$, as seen in Theorem B.1. Theorem B.1 also provides sufficient conditions (i.e. the non-extremality of the constraint levels) for the existence of Lagrange multipliers. This is non-trivial, since strong duality is not sufficient for the existence of Lagrange multipliers—[5, Example 5.3.3] provides an example where $\mathcal{M}_{0}$ is both convex and closed but $\text{Opt}(\text{D}_{0})=\emptyset$.
@@ B.1.5 Constraint perturbations
Upper bound . Let $\lambda\in\mathbb{R}_{+}^{I}$ and $\mu\in\mathbb{R}^{J}$ be arbitrarily chosen. Let $\mu_{+}$ and $\mu_{-}$ be defined pointwise as $\mu_{+,j}=\max\left\{0,\mu_{j}\right\}$ and $\mu_{-,j}=\max\left\{0,-\mu_{j}\right\}$. Clearly $\mu=\mu_{+}-\mu_{-}$ and $\left\|\mu_{+}\right\|_{1}+\left\|\mu_{-}\right\|_{1}=\left\|\mu\right\|_{1}$. Applying these facts, we can verify that $L_{0}$ can be rewritten as follows,
@@
Lower bound . Now let us prove the lower bound. Let $\lambda\in\mathbb{R}_{+}^{I}$ and $\mu_{+},\mu_{-}\in\mathbb{R}_{+}^{J}$ be arbitrarily chosen. Then we can rewrite $L_{\epsilon}$ as follows,
@@ B.2.1 Convexity of $\mathcal{M}_{\phi}$
Recall that $\mathcal{M}_{\phi}=F_{\phi}(\Phi)+\mathbb{R}_{+}^{1+I}\times\mathbf{0}^{J}$. Clearly $\mathbb{R}_{+}^{1+I}\times\mathbf{0}^{J}$ is a convex set, and since the Minkowski sum of two convex sets is convex ([50] Theorem 3.1), therefore $\mathcal{M}_{\phi}$ is also convex, which concludes the proof. ∎
@@ B.2.2 Strong duality of (Pϕ)
The premise of the proposition, along with atomlessness from Assumption 2 allows us to invoke Lemma B.5 to conclude that $\mathcal{M}_{\phi}$ is a convex set.
@@ B.3 Duality gap
Note that the set $\Theta_{\nu}$ (defined in Section 3) is nothing but the feasibility set of ($\text{P}_{\theta\nu}$), i.e. $\text{Feas}(\text{P}_{\theta\nu})$. Define the Lagrangian $L_{\theta\nu}(f_{\theta},\gamma)$ of ($\text{P}_{\theta\nu}$) similarly to (1), where the dual variable $\gamma\in\mathbb{R}_{+}^{I+2J}$ now corresponds to $I+2J$ inequality constraints. Let its dual value be $D^{\star}_{\theta\nu}$, defined similarly to $D^{\star}$ (Section 3).
We define its functional version,
@@
Let its Lagrangian and dual value be denoted as $L_{\phi\nu}(\phi,\gamma)$ and $D^{\star}_{\phi\nu}$ respectively, where $\gamma\in\mathbb{R}_{+}^{I+2J}$. We will bound the duality gap of (P) based on (5) which uses the functional strong duality relation $P^{\star}_{\phi}=D^{\star}_{\phi}$, and which we further decompose as,
@@ B.3.4 Duality gap of ($\text{P}_{\theta\nu}$)
Let $\gamma\in\mathbb{R}_{+}^{I+2J}$ denote the dual variable for ($\text{P}_{\theta\nu}$) or ($\text{P}_{\phi\nu}$). Clearly $L_{\theta\nu}(f_{\theta},\gamma)=L_{\phi\nu}(f_{\theta},\gamma)$. Therefore,
$\underset{\theta\in\Theta}{\text{inf}}\;L_{\theta\nu}(f_{\theta},\gamma)$ $=\underset{f_{\theta}\in\mathcal{H}}{\text{inf}}\;L_{\phi\nu}(f_{\theta},\gamma)\geq\underset{\phi\in\Phi}{\text{inf}}\;L_{\phi\nu}(\phi,\gamma)$
@@
$D^{\star}_{\theta\nu}=\sup_{\gamma\in\mathbb{R}_{+}^{I+2J}}\underset{\theta\in\Theta}{\text{inf}}\;L_{\theta\nu}(f_{\theta},\gamma)\geq\sup_{\gamma\in\mathbb{R}_{+}^{I+2J}}\underset{\phi\in\Phi}{\text{inf}}\;L_{\phi\nu}(\phi,\gamma)=D^{\star}_{\phi\nu}$
@@ B.4 Dual estimation error
$D^{\star}-\hat{D}^{\star}$ $=q(\lambda^{\star},\mu^{\star})-\hat{q}(\hat{\lambda}^{\star},\hat{\mu}^{\star})$
$\leq q(\lambda^{\star},\mu^{\star})-\hat{q}(\lambda^{\star},\mu^{\star}),$ (49)
where the inequality follows from the suboptimality of $(\lambda^{\star},\mu^{\star})$ with respect to ($\hat{\text{D}}$), i.e. because $\hat{q}(\lambda^{\star},\mu^{\star})\leq\hat{q}(\hat{\lambda}^{\star},\hat{\mu}^{\star})=\hat{D}^{\star}$. Let $\epsilon>0$ be arbitrary and let $\theta$ be $\epsilon$-optimal with respect to $\hat{q}(\lambda^{\star},\mu^{\star})$, i.e,
$\hat{q}(\lambda^{\star},\mu^{\star})\leq\hat{L}(f_{\theta},\lambda^{\star},\mu^{\star})\leq\hat{q}(\lambda^{\star},\mu^{\star})+\epsilon.$ (50)
Clearly, $q(\lambda^{\star},\mu^{\star})\leq L(f_{\theta},\lambda^{\star},\mu^{\star})$ since $q(\lambda^{\star},\mu^{\star})=\underset{\theta}{\text{inf}}\;L(f_{\theta},\lambda^{\star},\mu^{\star})$. This fact and (50), applied to (49), implies that,
$D^{\star}-\hat{D}^{\star}\leq q(\lambda^{\star},\mu^{\star})-\hat{q}(\lambda^{\star},\mu^{\star})\leq L(f_{\theta},\lambda^{\star},\mu^{\star})-\hat{L}(f_{\theta},\lambda^{\star},\mu^{\star})+\epsilon.$ (51)
Now let us expand the difference of Lagrangians in the RHS of (51),
@@
This time we pick some $\theta$ that is $\epsilon$-optimal with respect to $q(\hat{\lambda}^{\star},\hat{\mu}^{\star})$ (instead of $\hat{q}(\lambda^{\star},\mu^{\star})$). Since (53) holds for any $\theta$ (with probability greater than $1-(1+I+J)\delta$) therefore Equations (55) and (56) hold simultaneously with probability greater than the same bound, $1-(1+I+J)\delta$. Therefore, they together imply that $|D^{\star}-\hat{D}^{\star}|\leq(1+\max{}\left\{\left\|\gamma^{\star}\right\|_{1},\left\|\hat{\gamma}^{\star}\right\|_{1}\right\})\zeta^{\text{UC}}(N_{\text{min}},\delta)$, which completes the proof. ∎
@@
$\mathcal{M}=\left\{(f,\mathbf{u},\mathbf{v})\in\mathbb{R}^{1+I+J}\;\middle|\;\begin{aligned} \exists\theta\in\Theta\quad\text{ s.t.}\quad&\mathbb{E}_{\mathbb{D}}\left[\ell(f_{\theta}(\mathbf{x}),y)\right]=f,\\ &\mathbb{E}_{\mathbb{P}_{i}}\left[g_{i}(f_{\theta}(\mathbf{x}),y)\right]\leq u_{i},\quad\text{for }i=1,\ldots,I\text{,}\\ \text{ and }\quad&\mathbb{E}_{\mathbb{Q}_{j}}\left[h_{j}(f_{\theta}(\mathbf{x}),y)\right]=v_{j}\quad\text{for }j=1,\ldots,J\end{aligned}\right\}.$
@@
In this section we will provide guarantees for Algorithm 1, which approximately solves ($\hat{\text{D}}$). Recall the quantities related to the empirical dual problem, i.e. $\hat{L}(f_{\theta},\lambda,\mu),\hat{q}(\lambda,\mu)$ and $\hat{D}^{\star}$ (defined in Section 2). Also define the empirical constraint vector,
$\hat{C}(\theta)=\bigg[$ $\frac{1}{M_{1}}\sum_{m_{1}=1}^{M_{1}}g_{1}\big(f_{\theta}(\mathbf{x}_{m_{1}}),y_{m_{1}}\big)\big),\ldots,\frac{1}{M_{I}}\sum_{m_{I}=1}^{M_{I}}g_{I}\big(f_{\theta}(\mathbf{x}_{m_{I}}),y_{m_{I}}\big)\big),$
$\frac{1}{N_{1}}\sum_{n_{1}=1}^{N_{1}}h_{1}\big(f_{\theta}(\mathbf{x}_{n_{1}}),y_{n_{1}}\big),\ldots,\frac{1}{N_{J}}\sum_{n_{J}=1}^{N_{J}}h_{J}\big(f_{\theta}(\mathbf{x}_{n_{J}}),y_{n_{J}}\big)\bigg]^{\top}.$
Note that $\hat{C}$ has $I+J$ entries. Let $\gamma$ denote the tuple $(\lambda,\mu)$ in the sequel, and similarly for $\hat{\gamma}^{\star},\gamma^{(t-1)}$ etc. We will denote $\hat{q}(\lambda,\mu)$ as $\hat{q}(\gamma)$ and $\hat{L}(f_{\theta},\lambda,\mu)$ as $\hat{L}(f_{\theta},\gamma)$ for brevity.
We can prove that the set of supergradients for $\hat{q}(\gamma)$ at $\gamma$ is equal to the convex hull $\text{conv}\left(\left\{\hat{C}(\theta)\;|\;\theta\in\Theta,\hat{L}(f_{\theta},\gamma)=\hat{q}(\gamma)\right\}\right)$ (e.g. with [53, Theorem 2.87]). We will now show that approximate minimizers produce approximate supergradients.
@@
Let $\theta^{(t)}$ satisfy $\hat{q}(\gamma^{(t)})\leq\hat{L}(f_{\theta},\gamma^{(t)})\leq\hat{q}(\gamma^{(t)})+\rho$. Then, it follows that, for all $\gamma^{\prime}\in\mathbb{R}_{+}^{I}\times\mathbb{R}^{J}$
$\hat{q}(\gamma^{(t)})\geq\hat{q}(\gamma^{\prime})+\sum_{k=1}^{I+J}(\gamma^{(t)}_{k}-\gamma^{\prime}_{k})\hat{C}_{k}(\theta^{(t)})-\rho.$
@@
By the definition of the dual function, we have for any $\gamma^{\prime}\in\mathbb{R}_{+}^{I}\times\mathbb{R}^{J}$,